OpenStudy (anonymous):

6x+3y=3 3x=y what is x and y?

OpenStudy (jiteshmeghwal9):

3x=y x=y/3 \[2\cancel 6\left( y \over \cancel 3 \right)+3y=3\] \[2y+3y=3\]\[5y=3 \implies y={3 \over 5}\] \[\Huge{x={\frac{3}{5} \over 3}}\]\[x=5\]

OpenStudy (jiteshmeghwal9):

not sure

hartnn (hartnn):

x=1/5

hartnn (hartnn):

y is correct.

OpenStudy (jiteshmeghwal9):

ok